A large university is interested in the outcome of a course standardization process. They have taken a random sample of 100 student grades, across all instructors. The grades represent the proportion of problems answered correctly on a midterm exam. The sample proportion correct was calculated as 0.78.

  1. Construct a 90% confidence interval on the population proportion of correctly answered problems.

  2. Construct a 95% confidence interval on the population proportion of correctly answered problems.
Answer #1

a) Solution :-   

Given,

n = 100 ....... Sample size

   \hat p = 0.78 ....... sample proportion.

Our aim is to construct 90% confidence interval.

\therefore c = 0.90

\therefore\alpha = 1- c = 1- 0.90 = 0.10

\therefore  \alpha/2 = 0.10 \slash 2 = 0.05 and 1- \alpha /2 = 0.950

\therefore2012 = 1.645 (use z table)

Now , the margin of error is given by

E = 2012 *  Vp(1-P)/n

= 1.645 * \sqrt{} [ 0.78 *(1 - 0.78)/ 100]

= 0.0681

Now the confidence interval is given by

(\hat p - E)  < p < (\hat p + E)

( 0.78 - 0.0681 ) < p <   ( 0.78 + 0.0681 )

0.7119  < p < 0.8481  

Required 90% Confidence Interval is ( 0.7119 , 0.8481 )

b)  Solution :-   

Given,

n = 100 ....... Sample size

   \hat p = 0.78 ....... sample proportion.

Our aim is to construct 95% confidence interval.

\therefore c = 0.95

\therefore\alpha = 1- c = 1- 0.95 = 0.05

\therefore  \alpha/2 = 0.025 and 1- \alpha /2 = 0.975

Search the probability 0.975 in the Z table and see corresponding z value

\therefore2012 = 1.96

Now , the margin of error is given by

E = 2012 *  Vp(1-P)/n

= 1.96 * \sqrt{} [ 0.78 *(1 - 0.78)/ 100]

= 0.0812

Now the confidence interval is given by

(\hat p - E)  < p < (\hat p + E)

( 0.78 - 0.0812 ) < p <   ( 0.78 + 0.0812 )

0.6988  < p < 0.8612  

Required 95% Confidence Interval is ( 0.6988 , 0.8612 )


answered by: ANURANJAN SARSAM